Please help.
Solve the following system by substitution
x + y = 12
y = 2x
Thanks
JMS

Answer by checkley71(8403)

Y=2X NOW REPLACE Y WITH (2X & SOLVE FOR Y IN THE OTHER EQUATION.X+2X=12
3X=12
X=12/3
X=4 ANSWER.
Y=2*4
Y=8 ANSWER.
PROOF
4+8=12
12=12